OpenStudy (anonymous):

how do you simplify this fraction while fiinding the excluded value? -x/4x-8

OpenStudy (phi):

\[ \frac{-x}{4x-8} \] ? you can't simplify very much. Maybe factor out a 4 to get \[ \frac{-x}{4(x-2)} \] and multiply top and bottom by -1 to get \[ \frac{x}{4(2-x)} \] to find the excluded values of x, Exclude any x value that makes the denominator zero, because divide by zero is not allowed